(.4*X)+(.4*Y)+(.2*Z)=70 Solve for Z
Answer by Alan3354(69443)

(.4*X)+(.4*Y)+(.2*Z)=70 Solve for Z
----------
z can be determined only in terms of the x and y
2x + 2y + z = 350
z = 350 - 2x - 2y